OpenStudy (anonymous):

4^5/4^2 Simplify I wanna say it's 4^7 but I just have a feeling that's not right

OpenStudy (anonymous):

4^5/4^2=4^(5-2) = 4^3 Remember when dividing subtract powers.

OpenStudy (anonymous):

\[\frac{ a^m }{ a^n } = a ^{m - n}\]\[\frac{ 4^5 }{4^2 }= 4^{5-2}\]